If h(x)=h(x-1) + 4x and h(3) = -17, then find h(7). Please help!
Answer by richwmiller(17219)

You can get h(4) because you know h(4 -1) = h(3).
Then you work your way up.
h(3) = -17
h(4) = h(x-1) + 4x
= h(4 - 1) + 4*4
= h(3) + 16
= -17 + 16
= -1
h(5) = h(x-1) + 4x
= h(5 - 1) + 4*5
= h(4) + 20
= -1 + 20
= 19
h(6) = h(x-1) + 4x
= h(6 - 1) + 4*6
= h(5) + 24
= 19 + 24
= 43
h(7) = h(x-1) + 4x
= h(7 - 1) + 4*7
= h(6) + 28
= 43 + 28
= 71